OpenStudy (anonymous):

write the equation of the line passing throught the points (-3, -8) and (-6, 2) in slope-intercept form

OpenStudy (anonymous):

y=-10/3x-18

OpenStudy (anonymous):

how is it -18?

OpenStudy (anonymous):

y=mx+b is the equation of a line. in order to find the value for 'm', you can use the handy-dandy equation '(y[1] - y[2])/(x[1] - x[2])' 10/-3 is your value for m, in this case. Now that you know this, you can plug in one of the two pairs of co-ordinates to solve for 'b'. y=(-10/3)x+b -8=(-10/3)(-3)+b -8=10+b 0=10+8+b -b=18 b=-18. So, you would have 'y=-10/3x-18'

OpenStudy (anonymous):

is lx-7l the same thing as (x + 7)?

OpenStudy (anonymous):

and if you want put it in slope-intercept form. it could be y+8=-10/3(x+3) or y-2=-10/3(x+6)

OpenStudy (anonymous):

ok